OpenStudy (anonymous):

Pleeeeeeeeeease help :) Which ordered pairs in the form (x, y) are solutions to the equation (x + 6)y = 9? Choose all answers that are correct. A. (−5, 9) B. (−3, 3) C. (3, 1) D. (7, 9)

OpenStudy (anonymous):

A and B have a negative number before the 5 and 3 (-5,9) (-3,3)

OpenStudy (haichi):

It is A and B

OpenStudy (anonymous):

Can you explain

OpenStudy (haichi):

Jes

OpenStudy (gorv):

itzz A B and C

OpenStudy (gorv):

plug in all the values if both side comes equal thn itz a solution

OpenStudy (haichi):



OpenStudy (gorv):

(-5+6)*9=9 1*9=9 9=9

OpenStudy (gorv):

so itzz a solution

OpenStudy (anonymous):

Ohh, okay. Thanksss so much!!! :)

OpenStudy (haichi):

-3+6-=3 3*3=9

OpenStudy (haichi):

Heh......

OpenStudy (gorv):

similarly other will be

OpenStudy (haichi):

:|